In how many ways can 10 people take seats in a row of 24 fixed seats so that no two persons take consecutive seats.

Grade:11

1 Answers

Vikas TU
3 years ago
Vacant seat = 24-10 = 14 
 x0 can be greater than x0+x1+x2 +.....+x10 = 14
x0 can be greater than equal to 0 .
x0 + xy1 + y2 +..............+y9 + x10 + 9 = 14 
x0 + y1 + y2 + ....... +y9 + x10 = 5 
vacant seat which can be filled = 15c10 
No of ways no 2 consecutive seats are taken = 10! * 15C10